OpenStudy (anonymous):

Evaluate Plz. 6[5(3-9)-1]+2 _______________ 7(8-2)+4

OpenStudy (skullpatrol):

Have you learned what a "grouping symbol" is yet :-)

OpenStudy (anonymous):

nope lol

OpenStudy (skullpatrol):

A grouping symbol is a device used to enclose an expression that should be simplified before other operations are performed.

OpenStudy (skullpatrol):

Example parentheses, brackets, fraction bar.

OpenStudy (skullpatrol):

7(8-2)+4 = 7(6)+4 = 42+4 = 46

OpenStudy (skullpatrol):

When there are no grouping symbols simplify in the following order: 1. Do all multiplications and divisions in order from left to right. 2. Do all additions and subtraction in order from left to right.

OpenStudy (theopenstudyowl):

@ineedhelpnowplz